OpenStudy (anonymous):

The sum of the measures of the interior angles of a convex polygon is 2160°. How many sides does the polygon have? A. 12 B. 13 C. 14 D. 15

OpenStudy (anonymous):

@Flvs.net

OpenStudy (flvs.net):

Sum of interior angles of a polygon of n sides is equal to (n.180-360) given 180n-360=2160 180n=2160+360 hence n=13 sides

OpenStudy (flvs.net):

Wait no

OpenStudy (anonymous):

thanks. :)

OpenStudy (flvs.net):

Actually, N=14 2160+360=2520 2520/180=14